OpenStudy (anonymous):

30 = -4.4 square root x + 38

OpenStudy (anonymous):

\[30 = -4.4\sqrt{x} +38\]

OpenStudy (anonymous):

x=1.81818182

OpenStudy (anonymous):

1.81818182 same as lindsay!

OpenStudy (anonymous):

how do you get this answer ? :/

OpenStudy (anonymous):

30 = -4.4*sqrt(x) +38 30-38 = -4.4*sqrt(x) -8 = -4.4*sqrt(x) -8/-4.4 = sqrt(x) 1.818182 = sqrt(x) (1.818182)^2 = x 3.3058 = x

OpenStudy (anonymous):

Am also in support of klmoebes answer. by the way are you satisfied with the solution